# Short-segment speaker verification with pre-trained models and multi-resolution encoder

### v1

@@ Abstract
Speaker verification (SV) utilizing features obtained from models pre-trained via self-supervised learning has recently demonstrated impressive performances. However, these pre-trained models (PTMs) usually have a temporal resolution of 20 ms, which is lower than typical filterbank features. It may be problematic especially for short-segment SV with an input segment shorter than 2 s, in which we need to extract as much information as possible from the input with a limited length. Although there have been approaches to utilize multi-resolution features from the HuBERT models, the window shifts were 320, 640, and 1600 samples when the sampling rate was 16 kHz and thus only lower resolution features were considered. In this study, we propose an SV system which utilizes PTM features along with filterbank features and those from the multi-resolution time domain encoder with window shifts of 25, 50, 100, and 200 samples. Experimental results on the VoxCeleb dataset with various input lengths showed consistent improvements over systems with various combinations of input features.
@@ 1 Introduction
Recently, there has been growing interest in SV systems that utilize speech representations extracted from models pre-trained via self-supervised learning (SSL) on large-scale speech data [8, 9, 10, 11, 12, 13, 14, 15]. In particular, various fusion methods to combine layer-wise outputs from pre-trained models (PTMs) to produce the input for the SV models have been proposed [9, 10, 11, 12, 13]. While they exhibited potential of the PTM-based features, the temporal resolutions for these features were fixed to be the same as those for the corresponding PTMs, which is typically 20 ms. In [18], multiple temporal resolutions in HuBERT [19] have been proposed by pre-training three HuBERT models at different temporal resolutions and fusing their outputs using hierarchical or parallel approaches. MR-HuBERT [20] further extended the original HuBERT with two resolutions through a hierarchical Transformer, where predefined hidden representations were downsampled to incorporate both high- and low-resolution masked unit predictions, respectively. Although exploiting multiple temporal resolutions for HuBERT was shown to be beneficial for a broad range of speech processing tasks in these papers, the temporal resolutions they utilized corresponded to 20, 40, and 100 ms of frame shifts, which may be too low for the short-segment SV. In [11], the features from the PTMs fused with a fine-grained fusion module (FGFM) were combined with the FBank features with a 10 ms temporal resolution to produce excellent performances.

### v2

@@ Abstract
Speaker verification (SV) utilizing features obtained from models pre-trained via self-supervised learning has recently demonstrated impressive performances. However, these pre-trained models (PTMs) usually have a temporal resolution of 20 ms, which is lower than typical filterbank features. It may be problematic especially for short-segment SV with an input segment shorter than 2 s, in which we need to extract as much information as possible from the input with a limited length. Although there have been approaches to utilize multi-resolution features from the HuBERT models, the window shifts were 20, 40, and 100 ms when the sampling rate was 16 kHz and thus only lower resolution features were considered. In this study, we propose an SV system which utilizes PTM features along with filterbank features and those from the multi-resolution time domain encoder with window shifts of 1.56, 3.13, 6.25, and 12.5 ms. Experimental results on the VoxCeleb dataset with various input lengths showed consistent improvements over systems with various combinations of input features.
@@ 1 Introduction
Recently, there has been growing interest in SV systems that utilize speech representations extracted from models pre-trained via self-supervised learning (SSL) on large-scale speech data [26, 5, 16, 19, 24, 25, 23, 20]. In particular, various fusion methods to combine layer-wise outputs from pre-trained SSL models (PTMs) to produce the input for the SV models have been proposed [5, 16, 19, 24, 25]. While they exhibited potential of the PTM-based features, the temporal resolutions for these features were fixed to be the same as those for the corresponding PTMs, which is typically 20 ms. In [28], multiple temporal resolutions in HuBERT [14] have been proposed by pre-training three HuBERT models at different temporal resolutions and fusing their outputs using hierarchical or parallel approaches. MR-HuBERT [27] further extended the original HuBERT with two resolutions through a hierarchical Transformer, where predefined hidden representations were downsampled to incorporate both high- and low-resolution masked unit predictions, respectively. Although exploiting multiple temporal resolutions for HuBERT was shown to be beneficial for a broad range of speech processing tasks in these papers, the temporal resolutions they utilized corresponded to 20, 40, and 100 ms of frame shifts, which may be too low for the short-segment SV. In [19], the features from the PTMs fused with a fine-grained fusion module (FGFM) were combined with the FBank features with a 10 ms temporal resolution to produce excellent performances.